# Homological algebra of pro-Lie Polish abelian groups

### v1

@@ 3.3. Extensions of abelian Polish pro-Lie groups
In view of the results of [19, Chapter 3], in order to establish Theorem 3.14 it suffices to prove that the class of pro-Lie Polish abelian groups is closed under extensions.
@@
$\begin{array}[]{ccc}A&=&B\\ p_{n}\downarrow&&\downarrow\eta_{n}\\ A_{n}&\overset{\sigma_{n}}{\rightarrow}&B_{n}\end{array}$
@@
$\left(\left(t_{n},b_{n}\right)+\Xi_{n}\right)_{n\in\omega}=\left(\left(t_{n},b_{0}+a_{n}\right)+\Xi_{n}\right)_{n\in\omega}=\left(\left(t_{n}+s_{n},b_{0}\right)+\Xi_{n}\right)_{n\in\omega}=\left(\left(0,b+t+s\right)+\Xi_{n}\right)_{n\in\omega}=\eta\left(b_{0}+t+s\right)$
@@
Lemma 3.18.
Suppose that $A\rightarrow B\rightarrow C$ is a short exact sequence of abelian Polish groups, where $C\cong C_{0}\oplus C_{1}$. Then there exist closed subgroups $B_{0},B_{1}$ of $B$ such that $A\rightarrow B_{0}\oplus B_{1}\rightarrow C$ is a short exact sequence, and there exist short exact sequences $A\rightarrow B_{0}\rightarrow C_{0}$ and $A\rightarrow B_{1}\rightarrow C_{1}$.
Proof.
If $\pi:B\rightarrow C$ is the quotient map, identifying $C$ with $C_{0}\oplus C_{1}$ we can simply define $B_{i}:=\pi^{-1}\left(C_{i}\right)$ for $i\in\left\{0,1\right\}$. ∎
@@
We identify $C$ with $\prod_{n\in\omega}A_{n}$, and we identify $V$ with $\mathbb{R}$ and with a closed subgroup of $X$. For $n\in\omega$, we write $C_{>k}=\left\{x\in C:\forall i\leq k,x_{i}=0\right\}$. Fix a compatible complete invariant metric $d$ on $X$. For $x\in V$ we let $\left|x\right|$ be its absolute value in $\mathbb{R}$.
@@
We identify $\mathbb{T}$ with a closed subgroup of $X$ and with $\mathbb{R}/\mathbb{Z}$. We let $d_{\mathbb{R}}$ be metric on $\mathbb{T}$ defined by $d_{\mathbb{Z}}\left(t+\mathbb{Z},s+\mathbb{Z}\right)=\left|t-s\right|$ for $t,s\in[0,1)$. By [4, Proposition 4.6], we have a continuous right inverse $\varphi:C\rightarrow X$ for $\pi$. Define the continuous $\kappa:C\times C\rightarrow\mathbb{T}$ by $\kappa\left(x,y\right)=\varphi\left(x+y\right)-\varphi\left(x\right)-\varphi\left(y\right)$. Since $\kappa$ is continuous, there exists $n_{0}\in\omega$ such that $d_{\mathbb{Z}}\left(\kappa(x,y),0\right)<1/2$ for every $x\in C_{\geq n_{0}}:=\left\{x\in C:\forall i<n_{0},x_{i}=0\right\}$.
Since $\mathrm{Ext}_{\mathrm{Yon}}\left(C_{<n_{0}},\mathbb{T}\right)=0$ by injectivity of $\mathbb{T}$ in the category of locally compact abelian Polish groups, where $C_{<n_{0}}=\left\{x\in C:\forall i\geq n_{0},x_{i}=0\right\}$, it suffices to show that the function $\pi|_{\pi^{-1}\left(C_{\geq n}\right)}:\pi^{-1}\left(C_{\geq n}\right)\rightarrow C_{\geq n}$ has a right inverse that is a continuous group homomorphism. Thus, we can assume without loss of generality that $n=0$ and $d_{\mathbb{Z}}\left(\kappa\left(x,y\right),0\right)<1/2$ for every $x\in C$.
This implies that there exists a continuous $2$-cocycle $c:C\times C\rightarrow\mathbb{R}$ such that $c\left(x,y\right)+\mathbb{Z}=\kappa\left(x,y\right)$ for every $x,y\in C$. By Lemma 3.21 and Lemma 3.2, we have that $\mathrm{Ext}_{\mathrm{Yon}}\left(C,\mathbb{R}\right)=0$. Therefore, $c$ is a coboundary, and hence $\kappa$ is a coboundary as well. This implies that $\varphi$ has a right inverse that is a continuous homomorphism. ∎
@@
By Lemma 3.15 it suffices to consider the case when $L=\mathbb{R}$ or $L=\mathbb{T}$. In this case, Lemma 3.21 and Lemma 3.22 prove the result when $C$ is product of countable groups. If $C$ is an arbitrary non-Archimedean Polish abelian group, then by Lemma 3.23 there exists a surjective continuous homomorphism $\left(\mathbb{Z}^{\left(\omega\right)}\right)^{\omega}\rightarrow C$. Considering the pullback diagram
$\begin{array}[]{ccc}Y&\rightarrow&\left(\mathbb{Z}^{\left(\omega\right)}\right)^{\omega}\\ \downarrow&&\downarrow\\ X&\rightarrow&C\end{array}$
@@
whose rows and columns are short exact sequences. Then we have that $Y$ is pro-Lie by the above remarks, and hence $X$ is pro-Lie being quotient of a pro-Lie group by a closed subgroup. By Lemma 3.7, we conclude that the short exact sequence $L\rightarrow X\rightarrow C$ splits. ∎
@@
We just need to prove that if $A$ and $C$ are pro-Lie, then $B$ is pro-Lie, as the other implication is established in [19, Chapter 3]. By Lemma 3.9, we have a short exact sequence $D\rightarrow C\rightarrow T$ where $D$ is non-Archimedean and $T$ is a torus group. Considering the pullback diagram
@@ 3.4. Type decomposition for pro-Lie Polish abelian groups
We recall the type decomposition for locally compact Polish abelian groups as described by Hoffmann and Spitzweck in [21, Section 2].
@@
If $A$ is a pro-Lie Polish abelian group, then by [19, Theorem 5.20] we have that $A$ has a largest closed vector subgroup, which we denote by $A_{\mathbb{R}}$. Then we have that $A=A_{\mathbb{R}}\oplus B$ where $B$ is vector-free. We also let $A_{\mathbb{S}^{1}}$ be the connected component of the trivial element in $B$, which is compact [19, Theorem 24(ii)]. We let $c\left(A\right)$ be the connected component of the trivial element in $A$.
Lemma 3.29.
We have that $\mathbb{R}^{\omega}$ is projective in $\mathbf{proLiePAb}$.
Proof.
By [19, Theorem A2.12], $\mathrm{Ran}\left(\varphi|_{V}\right)$ is a closed subgroup and a topological direct summand of $\mathbb{R}^{\omega}$. Let $\left(N_{k}^{V}\right)$ be a cofinal sequence in $\mathcal{N}(V)$, $\left(N_{k}^{Y}\right)$ be a cofinal sequence in $\mathcal{N}(Y)$, and set $N_{k}:=N_{k}^{V}\oplus N_{k}^{Y}$ for $k\in\omega$. Thus, $\left(N_{k}\right)$ is a cofinal sequence in $\mathcal{N}\left(X\right)$. For $n\in\omega$, set
@@
We can write $A^{\left(k\right)}=V^{\left(k\right)}\oplus T^{\left(k\right)}$ where $V^{\left(k\right)}$ is a finite-dimensional vector group and $T^{\left(k\right)}$ is a finite-dimensional torus group. Furthermore, since $\mathrm{Hom}\left(T^{\left(k+1\right)},V^{\left(k\right)}\right)=0$ for every $k\in\omega$, $\left(V^{\left(k\right)}\right)$ is an epimorphic tower. Therefore, we have an exact sequence
Notice that, for every $k\in\omega$, if $d_{k}$ is the dimension of $T^{\left(k\right)}$, then $d_{k}+1$ is the maximum length of a chain of compact connected subgroups of $T^{\left(k\right)}$. It follows that the chain of subgroups $\left(\mathrm{Ran}\left(T^{\left(n\right)}\rightarrow T^{\left(k\right)}\right)\right)_{n>k}$ stabilizes. As this holds for every $k\in\omega$, $\mathrm{\mathrm{lim}}_{k}^{1}T^{\left(k\right)}=0$. By projectivity of vector groups, and since $\mathrm{Ran}\left(V^{\left(k+1\right)}\rightarrow V^{\left(k\right)}\right)$ is a closed vector subgroups of $V^{\left(k\right)}$ for every $k\in\omega$ by [19, Theorem A2.12], we have that $V:=\mathrm{\mathrm{lim}}_{k}V^{\left(k\right)}$ is a vector group. Hence, by projectivity of vector groups, we have that $\mathrm{\mathrm{lim}}_{k}A^{\left(k\right)}\cong\mathrm{\mathrm{lim}}_{k}T^{\left(k\right)}\oplus V$, where $\mathrm{\mathrm{lim}}_{k}T^{\left(k\right)}$ is connected by the previous lemma. ∎
@@
Let $A$ be a pro-Lie Polish abelian group. Suppose that $\left(A^{\left(k\right)}\right)$ is an epimorphic tower of Lie groups with $A=\mathrm{\mathrm{lim}}_{k}A^{\left(k\right)}$. Then $c\left(A\right)=\mathrm{\mathrm{lim}}_{k}c\left(A^{\left(k\right)}\right)$.
@@
We have that $\mathrm{\mathrm{lim}}_{k}c\left(A^{\left(k\right)}\right)$ is closed, and connected by Lemma 3.31. This shows that $\mathrm{\mathrm{lim}}_{k}c\left(A^{\left(k\right)}\right)\subseteq c\left(A\right)$.
By Lemma 3.33, $\left(c\left(A^{\left(k\right)}\right)\right)$ is an epimorphic tower. Let $\pi^{\left(n\right)}:\mathrm{\mathrm{lim}}_{k}A^{\left(k\right)}\rightarrow A^{\left(n\right)}$ be the canonical projection. Observe that $\pi^{\left(n\right)}\left(\mathrm{\mathrm{lim}}_{k}c(A^{\left(k\right)})\right)=c\left(A^{\left(n\right)}\right)$. If $U\subseteq\mathrm{\mathrm{lim}}_{k}c\left(A^{\left(k\right)}\right)$ is clopen, then $\pi^{\left(n\right)}\left(U\right)\subseteq c\left(A^{\left(n\right)}\right)$ is clopen, and hence $\pi^{\left(n\right)}\left(U\right)=c\left(A^{\left(n\right)}\right)$. This shows that $U$ is dense in $\mathrm{\mathrm{lim}}_{k}c\left(A^{\left(k\right)}\right)$, and hence equal to $\mathrm{\mathrm{lim}}_{k}c\left(A^{\left(k\right)}\right)$. ∎
@@
where $C^{\left(k\right)}$ is countable. Since $(c\left(A\right)^{\left(k\right)})$ is an epimorphic tower by Lemma 3.33, this implies by Lemma 3.34 that we have a short exact sequence
@@
Since $c\left(A\right)=\mathrm{\mathrm{lim}}_{k}c(A^{\left(k\right)})$, $A=\mathrm{\mathrm{lim}}_{k}A^{\left(k\right)}$, and $\mathrm{\mathrm{lim}}_{k}C^{\left(k\right)}$ is non-Archimedean, the conclusion follows. ∎
@@
For a locally compact Polish abelian group $A$, it is proved in [21, Proposition 2.2] that there exist canonical short exact sequences $F_{\mathbb{Z}}A\rightarrow A\rightarrow A_{\mathbb{Z}}$ and $A_{\mathbb{S}^{1}}\rightarrow F_{\mathbb{Z}}A\rightarrow A_{\mathbb{A}}$ where $A_{\mathbb{Z}}$, $A_{\mathbb{S}^{1}}$, and $A_{\mathbb{A}}$ have type $\mathbb{Z}$, $\mathbb{S}^{1}$, and $\mathbb{A}$, respectively. Furthermore, we have $A_{\mathbb{A}}=A_{\mathbb{R}}\oplus A_{\mathrm{t}}$ where $A_{\mathbb{R}}$ is a finite-dimensional vector group and $A_{\mathrm{t}}$ is a topological torsion group.
If $A$ is a pro-Lie Polish abelian group, we define $A_{\mathrm{t}}$ to be the largest topological torsion closed subgroup of $A\left/A_{0}\right.$, and $F_{\mathbb{Z}}A$ to be the preimage of $A_{\mathrm{t}}$ under the quotient map $A\rightarrow A\left/A_{0}\right.$. Define also $A_{\mathbb{Z}}:=A\left/F_{\mathbb{Z}}A\right.$ and $A_{\mathbb{S}^{1}}=\mathrm{comp}\left(A_{0}\right)$, which is the set of elements of $A_{0}$ that generate a subgroup with compact closure. Then we have that $A_{0}=A_{\mathbb{R}}\oplus A_{\mathbb{S}^{1}}$ where $A_{\mathbb{R}}$ is the largest closed vector subgroup of $A$. We also set $A_{\mathbb{A}}=A_{\mathbb{R}}\oplus A_{\mathrm{t}}$. Then we have canonical exact sequences
@@
This decomposition for pro-Lie Polish abelian groups can be reformulated in terms of torsion pars, showing that pro-Lie Polish abelian groups of type $\mathbb{A}$, type $\mathbb{S}^{1}$, and subtype $\mathbb{Z}$ form fully exact subcategories of the category $\mathbf{proLiePAb}$. This also follows from Lemma 2.4, together with the fact that locally compact pro-Lie Polish abelian groups of type $\mathbb{A}$, $\mathbb{S}^{1}$, and type $\mathbb{Z}$, respectively, form a fully exact subcategories of $\mathbf{LCPAb}$.

### v2

@@ 3.3. Extensions of abelian Polish pro-Lie groups
In view of the results of [20, Chapter 3], in order to establish Theorem 3.14 it suffices to prove that the class of pro-Lie Polish abelian groups is closed under extensions.
@@
$\begin{array}[]{ccc}A&\rightarrow&B\\ p_{n}\downarrow&&\downarrow\eta_{n}\\ A_{n}&\overset{\sigma_{n}}{\rightarrow}&B_{n}\end{array}$
@@
$\left(\left(t_{n},b_{n}\right)+\Xi_{n}\right)_{n\in\omega}=\left(\left(t_{n},b_{0}+a_{n}\right)+\Xi_{n}\right)_{n\in\omega}=\left(\left(t_{n}+s_{n},b_{0}\right)+\Xi_{n}\right)_{n\in\omega}=\left(\left(0,b_{0}+t+s\right)+\Xi_{n}\right)_{n\in\omega}=\eta\left(b_{0}+t+s\right)$
@@
We identify $C$ with $\prod_{n\in\omega}A_{n}$, and we identify $V$ with $\mathbb{R}$ and with a closed subgroup of $X$. For $n\in\omega$, we write $C_{>k}=\left\{x\in C:\forall i\leq k,x_{i}=0\right\}$. Fix a compatible complete invariant metric $d$ on $X$ [19, Exercise 2.2.9]. For $x\in V$ we let $\left|x\right|$ be its absolute value in $\mathbb{R}$.
@@
We identify $\mathbb{T}$ with a closed subgroup of $X$ and with $\mathbb{R}/\mathbb{Z}$. We let $d_{\mathbb{T}}$ be a compatible complete invariant metric on $\mathbb{T}$. For $\varepsilon>0$ define $\mathbb{T}_{\varepsilon}=\left\{x\in\mathbb{T}:d_{\mathbb{T}}\left(x,0\right)<\varepsilon\right\}$. Fix $\varepsilon>0$ such that there exists a continuous function $\rho:\mathbb{T}_{3\varepsilon}\rightarrow\mathbb{R}$ such that $\rho\left(z\right)+\mathbb{Z}=z$ for every $z\in\mathbb{T}_{3\varepsilon}$, and $\rho\left(x+y\right)=\rho\left(x\right)+\rho\left(y\right)$ for all $x,y\in\mathbb{T}_{\varepsilon}$.
By [4, Proposition 4.6], we have a continuous right inverse $\varphi:C\rightarrow X$ for $\pi$. Define the continuous $\kappa:C\times C\rightarrow\mathbb{T}$ by $\kappa\left(x,y\right)=\varphi\left(x+y\right)-\varphi\left(x\right)-\varphi\left(y\right)$. Since $\kappa$ is continuous, there exists $n_{0}\in\omega$ such that $d_{\mathbb{T}}\left(\kappa(x,y),0\right)<\varepsilon$ for every $x\in C_{\geq n_{0}}:=\left\{x\in C:\forall i<n_{0},x_{i}=0\right\}$.
Since $\mathrm{Ext}_{\mathrm{Yon}}\left(C_{<n_{0}},\mathbb{T}\right)=0$ by injectivity of $\mathbb{T}$ in the category of locally compact abelian Polish groups, where $C_{<n_{0}}=\left\{x\in C:\forall i\geq n_{0},x_{i}=0\right\}$, it suffices to show that the function $\pi|_{\pi^{-1}\left(C_{\geq n}\right)}:\pi^{-1}\left(C_{\geq n}\right)\rightarrow C_{\geq n}$ has a right inverse that is a continuous group homomorphism. Thus, we can assume without loss of generality that $n=0$ and $d_{\mathbb{T}}\left(\kappa\left(x,y\right),0\right)<\varepsilon$ for every $x\in C$.
This implies that, setting $c:=\rho\circ\kappa$, one obtains a continuous $2$-cocycle $c:C\times C\rightarrow\mathbb{R}$ such that $c\left(x,y\right)+\mathbb{Z}=\kappa\left(x,y\right)$ for every $x,y\in C$. By Lemma 3.20 and Lemma 3.2, we have that $\mathrm{Ext}_{\mathrm{Yon}}\left(C,\mathbb{R}\right)=0$. Therefore, $c$ is a coboundary, and hence $\kappa$ is a coboundary as well. This implies that $\varphi$ has a right inverse that is a continuous homomorphism. ∎
@@
By Lemma 3.15 it suffices to consider the case when $L=\mathbb{R}$ or $L=\mathbb{T}$. In this case, Lemma 3.20 and Lemma 3.21 prove the result when $C$ is product of countable groups. If $C$ is an arbitrary non-Archimedean Polish abelian group, then by Lemma 3.22 there exists a surjective continuous homomorphism $g:\left(\mathbb{Z}^{\left(\omega\right)}\right)^{\omega}\rightarrow C$. Considering the pullback diagram
$\begin{array}[]{ccc}Y&\rightarrow&\left(\mathbb{Z}^{\left(\omega\right)}\right)^{\omega}\\ \downarrow&&\downarrow g\\ X&\rightarrow&C\end{array}$
@@
whose rows and columns are short exact sequences. Then the exact sequence $L\rightarrow Y\rightarrow\left(\mathbb{Z}^{\left(\omega\right)}\right)^{\omega}$ splits as explained at the beginning of the proof of this lemma, hence $Y\cong L\oplus\left(\mathbb{Z}^{\left(\omega\right)}\right)^{\omega}$ is pro-Lie. Therefore, $X$ is pro-Lie, being quotient of a pro-Lie group by a closed subgroup. (Notice that a surjective continuous homomorphism between Polish groups is a quotient mapping by the Open Mapping Theorem for Polish groups; see [19, Corollary 2.3.4].) By Lemma 3.7, we conclude that the short exact sequence $L\rightarrow X\rightarrow C$ splits. ∎
@@
We just need to prove that if $A$ and $C$ are pro-Lie, then $B$ is pro-Lie, as the other implication is established in [20, Chapter 3]. By Lemma 3.9, we have a short exact sequence $D\rightarrow C\rightarrow T$ where $D$ is non-Archimedean and $T$ is a torus group. Considering the pullback diagram
@@ 3.4. Type decomposition for pro-Lie Polish abelian groups
We recall the type decomposition for locally compact Polish abelian groups as described by Hoffmann and Spitzweck in [22, Section 2].
@@
For a locally compact Polish abelian group $A$, it is proved in [22, Proposition 2.2] that there exist canonical short exact sequences $F_{\mathbb{Z}}A\rightarrow A\rightarrow A_{\mathbb{Z}}$ and $A_{\mathbb{S}^{1}}\rightarrow F_{\mathbb{Z}}A\rightarrow A_{\mathbb{A}}$ where $A_{\mathbb{Z}}$, $A_{\mathbb{S}^{1}}$, and $A_{\mathbb{A}}$ have type $\mathbb{Z}$, $\mathbb{S}^{1}$, and $\mathbb{A}$, respectively. Furthermore, we have $A_{\mathbb{A}}=A_{\mathbb{R}}\oplus A_{\mathrm{t}}$ where $A_{\mathbb{R}}$ is a finite-dimensional vector group and $A_{\mathrm{t}}$ is a topological torsion group.
@@
If $A$ is a pro-Lie Polish abelian group, then by [20, Theorem 5.20] we have that $A$ has a largest closed vector subgroup, which we denote by $A_{\mathbb{R}}$. Then we have that $A=A_{\mathbb{R}}\oplus B$ where $B$ is vector-free. We also let $A_{\mathbb{S}^{1}}$ be the connected component of the trivial element in $B$, which is compact [20, Theorem 24(ii)]. We let $c\left(A\right)$ be the connected component of the trivial element in $A$.
@@
If $V$ is a locally compact vector group, then $V$ is projective in the category of pro-Lie Polish abelian groups.x
@@
Suppose that $Y$ is a pro-Lie Polish abelian group and $\varphi:Y\rightarrow V$ is a surjective continuous group homomorphism. Since $V$ is a Lie group, we can find a tower $\left(Y_{n},\eta_{n}:Y_{n+1}\rightarrow Y_{n}\right)$ of abelian Lie groups and a continuous group homomorphism $\varphi_{0}:Y_{0}\rightarrow V$ such that $Y=\mathrm{\mathrm{lim}}_{n}Y_{n}$ and $\varphi=\varphi_{0}\circ\pi_{0}$ where $\pi_{0}:Y\rightarrow Y_{0}$ is the canonical map. By projectivity of $V$ in the category of locally compact Polish abelian groups, we can find a continuous homomorphism $\psi_{0}:V\rightarrow Y_{0}$ that is a right inverse for $\varphi_{0}$. Since the image of $\psi_{0}$ is a closed subgroup of $Y_{0}$ isomorphic to $V$, we can find a continuous homomorphism $\psi_{1}:\psi_{0}\left(V\right)\rightarrow Y_{1}$ that is a right inverse for $\eta_{0}|_{\eta_{0}^{-1}(\psi_{0}(V))}$, whence $\psi_{1}\circ\psi_{0}$ is a continuous homomorphism that is a right inverse for $\varphi_{0}\circ\eta_{0}$. Proceeding in this fashion, we can define recursively a continuous homomorphisms $\psi_{n}:\left(\psi_{n-1}\circ\cdots\circ\psi_{0}\right)\left(V\right)\rightarrow Y_{n}$ that is a right inverse for $\eta_{n}|_{\eta_{n}^{-1}((\psi_{n}\circ\cdots\circ\psi_{0})(V))}$, whence $\psi_{n}\circ\cdots\circ\psi_{0}$ is a continuous homomorphism that is a right inverse for $\varphi_{0}\circ\eta_{0}\circ\cdots\circ\eta_{n-1}$.
One can then define the continuous group homomorphism $\psi:V\rightarrow Y$, $x\mapsto\left(\left(\psi_{n}\circ\cdots\circ\psi_{0}\right)(x)\right)_{n\in\omega}$, which is a right inverse for $\varphi$. ∎
Corollary 3.30.
If $Y$ is a vector-free pro-Lie Polish abelian group, then for every $N\in\mathcal{N}\left(Y\right)$, $Y/N$ is a vector-free abelian Lie group.
@@
Let $A$ be a vector-free pro-Lie Polish abelian group. Suppose that $\left(A^{\left(k\right)}\right)$ is an epimorphic tower of Lie groups with $A=\mathrm{\mathrm{lim}}_{k}A^{\left(k\right)}$. Then $c\left(A\right)=\mathrm{\mathrm{lim}}_{k}c\left(A^{\left(k\right)}\right)$ is a compact connected subgroups of $A$.
@@
Notice that, for every $k\in\omega$, $A^{\left(k\right)}$ is vector-free. We have that $\mathrm{\mathrm{lim}}_{k}c(A^{\left(k\right)})$ is compact, and connected by Lemma 3.28. This shows that $\mathrm{\mathrm{lim}}_{k}c(A^{\left(k\right)})\subseteq c\left(A\right)$. The converse inclusion follow from functoriality of $c\left(A\right)$. ∎
@@
Suppose that $A$ is a vector-free pro-Lie Polish abelian group. Then $A\left/c\left(A\right)\right.$ is a non-Archimedean Polish abelian group.
@@
Let $\left(A^{\left(k\right)}\right)$ be an epimorphic tower of (necessarily vector-free) Lie Polish abelian groups with $A=\mathrm{\mathrm{lim}}_{k}A^{\left(k\right)}$. For every $k\in\omega$, we have a short exact sequence
$c(A^{\left(k\right)})\rightarrow A^{\left(k\right)}\rightarrow C^{\left(k\right)}$
where $C^{\left(k\right)}$ is countable. Since $(c\left(A\right)^{\left(k\right)})$ is an epimorphic tower by Lemma 3.32, this implies that we have a short exact sequence
@@
Since $c\left(A\right)=\mathrm{\mathrm{lim}}_{k}c(A^{\left(k\right)})$ by Lemma 3.32, $A=\mathrm{\mathrm{lim}}_{k}A^{\left(k\right)}$, and $\mathrm{\mathrm{lim}}_{k}C^{\left(k\right)}$ is non-Archimedean, the conclusion follows. ∎
@@
Lemma 3.36.
Proof.
As in the proof of Lemma 3.28, we can assume without loss of generality that $\left(A^{\left(k\right)}\right)$ is an epimorphic tower. We can write $A^{\left(k\right)}=V^{\left(k\right)}\oplus T^{\left(k\right)}$ where $V^{\left(k\right)}$ is a finite-dimensional vector group and $T^{\left(k\right)}$ is a finite-dimensional torus group. Furthermore, since $\mathrm{Hom}\left(T^{\left(k+1\right)},V^{\left(k\right)}\right)=0$ for every $k\in\omega$, $\left(V^{\left(k\right)}\right)$ is an epimorphic tower. Therefore, we have an exact sequence
Notice that, for every $k\in\omega$, if $d_{k}$ is the dimension of $T^{\left(k\right)}$, then $d_{k}+1$ is the maximum length of a chain of compact connected subgroups of $T^{\left(k\right)}$. It follows that the chain of subgroups $\left(\mathrm{Ran}\left(T^{\left(n\right)}\rightarrow T^{\left(k\right)}\right)\right)_{n>k}$ stabilizes. As this holds for every $k\in\omega$, $\mathrm{\mathrm{lim}}_{k}^{1}T^{\left(k\right)}=0$. By projectivity of vector groups, and since $\mathrm{Ran}\left(V^{\left(k+1\right)}\rightarrow V^{\left(k\right)}\right)$ is a closed vector subgroup of $V^{\left(k\right)}$ for every $k\in\omega$ by [20, Theorem A2.12], we have that $V:=\mathrm{\mathrm{lim}}_{k}V^{\left(k\right)}$ is a vector group. Hence, by projectivity of vector groups, we have that $\mathrm{\mathrm{lim}}_{k}A^{\left(k\right)}\cong\mathrm{\mathrm{lim}}_{k}T^{\left(k\right)}\oplus V$, where $\mathrm{\mathrm{lim}}_{k}T^{\left(k\right)}$ is connected by the previous lemma. ∎
Corollary 3.37.
Let $A$ be a pro-Lie Polish abelian group.
  1. (1)
    Suppose that $\left(A^{\left(k\right)}\right)$ is an epimorphic tower of Lie groups with $A=\mathrm{\mathrm{lim}}_{k}A^{\left(k\right)}$. Then $c\left(A\right)=\mathrm{\mathrm{lim}}_{k}c\left(A^{\left(k\right)}\right)$.
  2. (2)
    $A/c\left(A\right)$ is a non-Archimedean Polish abelian group.
Proof.
The proofs of Lemma 3.33 and Lemma 3.34 apply verbatim with Lemma 3.28 replaced with Lemma 3.36. ∎
Lemma 3.38.
Proof.
where $C^{\left(k\right)}$ is countable. Since $(c\left(A\right)^{\left(k\right)})$ is an epimorphic tower by Lemma 3.32, we have a short exact sequence
By Corollary 3.37, $c\left(A\right)=\mathrm{\mathrm{lim}}_{k}c(A^{\left(k\right)})$. Hence, $A/c\left(A\right)\cong\mathrm{\mathrm{lim}}_{k}C^{\left(k\right)}$ is non-Archimedean. ∎
Lemma 3.39.
Proof.
Lemma 3.40.
We have that $\mathbb{R}^{\omega}$ is projective in $\mathbf{proLiePAb}$.
Proof.
By [20, Theorem A2.12], $\mathrm{Ran}\left(\varphi|_{V}\right)$ is a closed subgroup and a topological direct summand of $\mathbb{R}^{\omega}$.
Let $\left(N_{k}^{V}\right)$ be a cofinal sequence in $\mathcal{N}(V)$, $\left(N_{k}^{Y}\right)$ be a cofinal sequence in $\mathcal{N}(Y)$, and set $N_{k}:=N_{k}^{V}\oplus N_{k}^{Y}$ for $k\in\omega$. Thus, $\left(N_{k}\right)$ is a cofinal sequence in $\mathcal{N}\left(X\right)$. For $n\in\omega$, set
If $A$ is a pro-Lie Polish abelian group, we define $A_{\mathrm{t}}$ to be the largest topological torsion closed subgroup of $A\left/c\left(A\right)\right.$, and $F_{\mathbb{Z}}A$ to be the preimage of $A_{\mathrm{t}}$ under the quotient map $A\rightarrow A\left/c\left(A\right)\right.$. Define also $A_{\mathbb{Z}}:=A\left/F_{\mathbb{Z}}A\right.$ and $A_{\mathbb{S}^{1}}=\mathrm{comp}\left(c\left(A\right)\right)$, which is the set of elements of $c\left(A\right)$ that generate a subgroup with compact closure.
By Lemma LABEL:Lemma:R-projective, we have that $c\left(A\right)=A_{\mathbb{R}}\oplus A_{\mathbb{S}^{1}}$ where $A_{\mathbb{R}}$ is the largest closed vector subgroup of $A$. We also set $A_{\mathbb{A}}=A_{\mathbb{R}}\oplus A_{\mathrm{t}}$. Then we have canonical exact sequences
@@
This decomposition for pro-Lie Polish abelian groups can be reformulated in terms of torsion pars, showing that pro-Lie Polish abelian groups of type $\mathbb{A}$, type $\mathbb{S}^{1}$, and subtype $\mathbb{Z}$ form fully exact subcategories of the category $\mathbf{proLiePAb}$. This also follows from Lemma LABEL:Lemma:thick-pro-category, together with the fact that locally compact pro-Lie Polish abelian groups of type $\mathbb{A}$, $\mathbb{S}^{1}$, and type $\mathbb{Z}$, respectively, form a fully exact subcategories of $\mathbf{LCPAb}$.